OpenStudy (anonymous):

if you are starting a salaried job at $22,000 per years and after 6 months on the job you get a 10 % raise what will be the total income for the year, and what will be the annual salary income at the end of the year?

OpenStudy (radar):

For the first 6 months (1/2 year) you will receive $11,000 The remaining 6 months will include the 10% raise which works out to be $11,000 + .10*$11,000=(11,000+1100)=$12,100 the year total salary is $23,100, His annual income at the end of the year is ($22,000+.1*22,000)=22,000+2200=$24,200

OpenStudy (anonymous):

thanks